what is the equation of a line that passes through
(-3,0) and (0,3)

Answer by Nate(3500)

slope is the difference of the y's divide by the difference of the x's
m=(0-3)/(-3-0)
m=(-3)/(-3)
m=1
the equation for finding the equation is point slope
y - first(y)=m(x - first(x))
y - 0=1(x - 3)
y=x-3 is your equation!